OpenStudy (anonymous):

Given the reaction 4NH3(g)+3O2(g)→2N2(g)+6H2O(l). What mass in grams of water is produced from 120.0 grams of NH3 reacting with excess oxygen.?

OpenStudy (aaronq):

You need to convert the mass of NH3 to moles first \( \sf moles=\dfrac{mass}{Molar~mass}\) The use the stoichiometric coefficients from the balanced equation to relate the moles of reactant to moles of product. \(\sf \dfrac{moles~of~NH_3}{NH_3's~coefficient}=\dfrac{moles~of~water}{water's~coefficient}\) find moles of water from the equation above by pluggin in what you know (everything except moles of water) Convert the moles of water to mass using the first equation.